### v1

@@ 1 Introduction
Thus, by the (ZFC) results of Section 4, the poset ${\mathbb{P}}(\omega^{\delta})$ either preserves $\omega_{1}$ and forces $|{\mathfrak{c}}|=|{\mathfrak{h}}|$, or collapses $\omega_{2}$ to $\omega$. So, it is natural to ask whether ${\mathbb{P}}(\omega^{\delta})$ collapses more cardinals (which, of course, depends on $|\delta|$, $\mathop{\rm cc}\nolimits({\mathbb{P}}(\omega^{\delta}))$, etc.) and a more ambitious task is to determine $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))$ (using, for example, some known characterizations of collapsing algebras). Here we note that, generally, for a poset ${\mathbb{P}}$ we have
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\cong\mathop{\rm Col}\nolimits(\lambda,\kappa)\Rightarrow\mathop{\rm Col}\nolimits(\lambda,\kappa)\hookrightarrow_{c}\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\Rightarrow 1_{\mathbb{P}}\Vdash|\check{\kappa}|=|\check{\lambda}|$
and by Kripke’s theorem (see [10], p. 227) for a Boolean algebra ${\mathbb{B}}$ of size $\leq\kappa$ we have $\mathop{\rm ro}\nolimits({\mathbb{B}}^{+}\times\mathop{\rm Col}\nolimits(\omega,\kappa)^{+})\cong\mathop{\rm Col}\nolimits(\omega,\kappa)$; thus the first iteration is “absorbed” by $\mathop{\rm Col}\nolimits(\omega,\kappa)$. Since $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|=|P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}|\leq 2^{|\delta|}$, the extreme case is when ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega_{1}$, in cases (A) and (B), and, in other cases, collapses $2^{|\delta|}$ to $\omega$. A more ambitious desire is to have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits(P(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{k},2^{|\delta|})$, for $k=1$ in the first case and for $k=0$ in the second. So, in Section 5 we consider that phenomenon of “maximal collapse”. For example, concerning Case (A), we show that if ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$, then ${\mathfrak{h}}(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))=\omega_{1}$ (in ZFC) and that, under some additional conditions (for example, if ${\mathfrak{h}}=\omega_{1}$ or if $0^{\sharp}$ does not exist and ${\mathfrak{b}}={\mathfrak{d}}$, e.g. if MA is true), we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits(P(\omega^{\delta+1})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. Regarding Cases (D) and (E) we show that $|\delta|=\mathop{\rm cf}\nolimits(\delta)=\kappa>\omega$ and $2^{\kappa}=\kappa^{+}$ implies that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$. A simple example shows that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$ is consistent. In Section 6 on the basis of the results from Sections 4 and 5 we analyze the general case, when $\alpha\geq\omega_{1}$.
@@ 2 Preliminaries
If ${\mathbb{P}}=\langle P,\leq\rangle$ is a preorder, the elements $p$ and $q$ of $P$ are incompatible, we write $p\perp q$, iff there is no $r\in P$ such that $r\leq p,q$. A set $A\subset P$ is an antichain iff $p\perp q$, for different $p,q\in A$. An element $p$ of $P$ is an atom iff each two elements $q,r\leq p$ are compatible. ${\mathbb{P}}$ is called: atomless iff it has no atoms; homogeneous iff it has a largest element and ${\mathbb{P}}\cong p\!\downarrow$, for each $p\in{\mathbb{P}}$, where $p\!\downarrow=\{q\in{\mathbb{P}}:q\leq p\}$. A set $D\subset P$ is called dense iff for each $p\in P$ there is $q\in D$ such that $q\leq p$; $D$ is called open iff $q\leq p\in D$ implies $q\in D$. A set $\Phi\subset P$ is a filter in ${\mathbb{P}}$ iff $\Phi\ni p\leq q$ implies $q\in\Phi$ and for each $p,q\in\Phi$ there is $r\in\Phi$ such that $r\leq p,q$. If $\kappa$ is a cardinal, ${\mathbb{P}}$ is called $\kappa$-closed iff for each $\gamma<\kappa$ each sequence $\langle p_{\alpha}:\alpha<\gamma\rangle$ in $P$, such that $\alpha<\beta\Rightarrow p_{\beta}\leq p_{\alpha}$, has a lower bound in $P$. $\omega_{1}$-closed preorders are called $\sigma$-closed. By ${\mathfrak{h}}({\mathbb{P}})$ we denote the distributivity number of a poset ${\mathbb{P}}$.
@@
If ${\mathbb{P}}_{i}$, $i\in I$, are $\kappa$-closed preorders, then $\prod_{i\in I}{\mathbb{P}}_{i}$ is $\kappa$-closed.
A partial order ${\mathbb{P}}=\langle P,\leq\rangle$ is called separative iff for each $p,q\in P$ satisfying $p\not\leq q$ there is $r\leq p$ such that $r\perp q$. The separative modification of ${\mathbb{P}}$ is the separative preorder $\mathop{\rm sm}\nolimits({\mathbb{P}})=\langle P,\leq^{*}\rangle$, where
$p\leq^{*}q\Leftrightarrow\forall r\leq p\;\exists s\leq r\;s\leq q.$ (8)
The separative quotient of ${\mathbb{P}}$ is the separative partial order $\mathop{\rm sq}\nolimits({\mathbb{P}})=\langle P/\!\!=^{*},\trianglelefteq\rangle$, where $p=^{*}q\Leftrightarrow p\leq^{*}q\land q\leq^{*}p\;$ and $\;[p]\trianglelefteq[q]\Leftrightarrow p\leq^{*}q$. Two preorders ${\mathbb{P}}$ and ${\mathbb{Q}}$ are called forcing equivalent iff they produce the same generic extensions.
Fact 2.2
Let ${\mathbb{P}},{\mathbb{Q}}$ and ${\mathbb{P}}_{i}$, $i\in I$, be partial orderings. Then
(a) ${\mathbb{P}}$, $\mathop{\rm sm}\nolimits({\mathbb{P}})$ and $\mathop{\rm sq}\nolimits({\mathbb{P}})$ are forcing equivalent forcing notions;
(b) $\mathop{\rm sm}\nolimits({\mathbb{P}})$ is $\kappa$-closed iff $\mathop{\rm sq}\nolimits({\mathbb{P}})$ is $\kappa$-closed;
(c) If $p_{0},p_{1},\dots p_{n}\in{\mathbb{P}}$, where $p_{n}\leq^{*}p_{n-1}\leq^{*}\dots\leq^{*}p_{0}$, then there is $q\in{\mathbb{P}}$ such that $q\leq p_{k}$, for all $k\leq n$.
(d) ${\mathbb{P}}\cong{\mathbb{Q}}$ implies that $\mathop{\rm sm}\nolimits{\mathbb{P}}\cong\mathop{\rm sm}\nolimits{\mathbb{Q}}$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}\cong\mathop{\rm sq}\nolimits{\mathbb{Q}}$;
(e) $\mathop{\rm sm}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})=\prod_{i\in I}\mathop{\rm sm}\nolimits{\mathbb{P}}_{i}$ and $\mathop{\rm sq}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})\cong\prod_{i\in I}\mathop{\rm sq}\nolimits{\mathbb{P}}_{i}$.
(f) If $X$ is an infinite set, ${\mathcal{I}}\subset P(X)$ an ideal containing $[X]^{<\omega}$ and ${\mathcal{I}}^{+}=P(X)\setminus{\mathcal{I}}$ the corresponding family of ${\mathcal{I}}$-positive sets, then sm+,normal-$\mathop{\rm sm}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle$ $=\langle{\mathcal{I}}^{+},\subset_{\mathcal{I}}\rangle$, where $A\subset_{\mathcal{I}}B\Leftrightarrow A\setminus B\in{\mathcal{I}}$, for $A,B\in{\mathcal{I}}^{+}$. Also $\mathop{\rm sq}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle=(P(X)/{\mathcal{I}})^{+}$.
(g) $A\not\perp_{\mathbb{P}}B$ iff $A\cap B\in{\mathcal{I}}^{+}$.
Proof. All the statements are folklore except, maybe, (c). For a proof of (c), by recursion we define the sequence $\langle q_{k}:k\leq m\rangle$ such that (i) $q_{0}=p_{n}$ and (ii) $q_{k}\leq q_{k-1},p_{n-k}$, for $0<k\leq n$. Then $q_{n}\leq p_{k}$, for all $k\leq n$. $\Box$
We recall that, if $\langle{\mathbb{P}}\leq_{\mathbb{P}},1_{\mathbb{P}}\rangle$ and $\langle{\mathbb{Q}}\leq_{\mathbb{Q}},1_{\mathbb{Q}}\rangle$ are preorders, then a mapping $f:{\mathbb{P}}\rightarrow{\mathbb{Q}}$ is a complete embedding, in notation $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$ iff
@@
If $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$, then ${\mathbb{Q}}$ is forcing equivalent to the two step iteration ${\mathbb{P}}\ast\langle\pi,\leq_{\pi},\check{1_{\mathbb{Q}}}\rangle$, where $1_{\mathbb{P}}\Vdash_{\mathbb{P}}\pi\subset\check{{\mathbb{Q}}}$ and for each $p\in{\mathbb{P}}$ and $q,q_{1},q_{2}\in{\mathbb{Q}}$
(a) $p\Vdash\check{q}\in\pi$ iff $p\in\mathop{\rm red}\nolimits(q)$;
(b) $p\Vdash\check{q_{1}}\leq_{\pi}\check{q_{2}}$ iff $q_{1}\leq_{\mathbb{Q}}q_{2}$ and $p\in\mathop{\rm red}\nolimits(q_{1})$.
@@
If $V\models\mathrm{GCH}$ and, in $V$, $\kappa>\omega$ is a regular cardinal and ${\mathbb{P}}=\mathop{\mathrm{Fn}}\nolimits(\kappa,2)$, then in $V[G]$ we have ${\mathfrak{c}}=\kappa$. If, in $V$, $\mu$ and $\theta$ are infinite cardinals, then (see [11], p. 246)
$(\theta^{\mu})^{V[G]}=(\max(\kappa,\theta)^{\mu})^{V}$ (9)
@@
Let $\kappa\geq 2$ and $\lambda\geq\omega$ be cardinals. The collapsing algebra $\mathop{\rm Col}\nolimits(\lambda,\kappa)$ is the Boolean completion $\mathop{\rm ro}\nolimits({}^{<\lambda}\kappa)$ of the reversed tree $\langle{}^{<\lambda}\kappa,\supset\rangle$. A Boolean algebra ${\mathbb{B}}$ is $(\lambda,\cdot,\kappa)$-nowhere distributive iff there are maximal antichains $A_{\alpha}$, $\alpha<\lambda$, such that each $b\in B^{+}$ is compatible with $\geq\kappa$ elements of some $A_{\alpha}$. (Then $\mathop{\rm cc}\nolimits({\mathbb{B}})>\kappa$.) Clearly, the algebra $\mathop{\rm Col}\nolimits(\lambda,\kappa)$ is $(\lambda,\cdot,\kappa)$-nowhere distributive and we have the following characterization.
Fact 2.5 (Balcar, Simon, [3], p. 338)
Let ${\mathbb{B}}$ be a homogeneous Boolean algebra, $\lambda\geq\omega$ a regular cardinal and $\kappa\geq 2$ a cardinal. If ${\mathbb{B}}$ is $(\lambda,\cdot,\kappa)$-nowhere distributive and contains a $\lambda$-closed dense subset, then $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,\kappa)\Leftrightarrow\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,\kappa^{<\lambda})\Leftrightarrow\pi({\mathbb{B}})=\kappa^{<\lambda}$.
@@
(Folklore) Let ${\mathbb{P}}$ be a separative atomless preorder of size $\kappa$ and $\lambda$ a regular cardinal. If ${\mathbb{P}}$ is $\lambda$-closed, $1_{\mathbb{P}}\Vdash|\check{\kappa}|=\check{\lambda}$ and $\kappa^{<\lambda}=\kappa$, then ${\mathbb{P}}\equiv_{forc}\mathop{\rm Col}\nolimits(\lambda,\kappa)$. In particular, if $1_{\mathbb{P}}\Vdash|\check{\kappa}|=\check{\omega}$, then ${\mathbb{P}}\equiv_{forc}\mathop{\rm Col}\nolimits(\omega,\kappa)$.
Let $\kappa\geq\omega$ be a cardinal and $\subset^{*}$ the preorder on the set $[\kappa]^{\kappa}$ defined by $A\subset^{*}B$ iff $|A\setminus B|<\kappa$. By ${\mathcal{P}}_{\kappa}$ we denote the quotient algebra $P(\kappa)/[\kappa]^{<\kappa}$.
@@
If $\kappa$ is an infinite cardinal, then
(a) The algebra ${\mathcal{P}}_{\kappa}$ is atomless, homogeneous, of size $2^{\kappa}$ and $\kappa^{++}\leq\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})\leq(2^{\kappa})^{+}$ (see [3], p. 372);
(b) (Balcar, Simon, [2]) If $\kappa>\mathop{\rm cf}\nolimits(\kappa)>\omega$, then the algebra ${\mathcal{P}}_{\kappa}$ is $(\omega,\cdot,\kappa^{+})$-nowhere distributive.
@@
If $\kappa$ is an infinite cardinal, then
(a) $\mathop{\rm sm}\nolimits\langle[\kappa]^{\kappa},\subset\rangle=\langle[\kappa]^{\kappa},\subset^{*}\rangle$ and $\mathop{\rm sq}\nolimits\langle[\kappa]^{\kappa},\subset\rangle={\mathcal{P}}_{\kappa}^{+}$;
(b) (Shelah, [23]) If $\kappa>2^{\mathop{\rm cf}\nolimits(\kappa)}>\mathop{\rm cf}\nolimits(\kappa)>\omega$, then $\mathop{\rm Col}\nolimits(\omega,\kappa^{+})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$.
(c) (Shelah, [23]) If $\kappa=\mathop{\rm cf}\nolimits(\kappa)>\omega$, then forcing by ${\mathcal{P}}_{\kappa}^{+}$ collapses each $\lambda<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ to $\omega$;
(d) (Baumgartner, [5]) It is consistent with ZFC that $2^{\omega_{1}}$ is arbitrarily large and that $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}$;
(e) (Balcar, Pelant and Simon, [1]) Forcing by ${\mathcal{P}}_{\omega}^{+}$ collapses $2^{\omega}$ to ${\mathfrak{h}}$;
(f) (Kojman, Shelah [9]) If $\kappa>\mathop{\rm cf}\nolimits(\kappa)=\omega$, then $\mathop{\rm Col}\nolimits(\omega_{1},\kappa^{\omega})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$ (see also [4]).
Fact 2.9
Let $\kappa\in\mathop{\mathrm{Card}}\nolimits$, $2^{\kappa}=\kappa^{+}$ and $2^{\mathop{\rm cf}\nolimits(\kappa)}=\mathop{\rm cf}\nolimits(\kappa)^{+}$. Then (see [3], p. 380)
(a) If $\mathop{\rm cf}\nolimits(\kappa)>\omega$, then $\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})\cong\mathop{\rm Col}\nolimits(\omega\,,2^{\kappa})$;
(b) If $\mathop{\rm cf}\nolimits(\kappa)=\omega$, then $\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\kappa})$.
Linear orders. Indecomposable ordinals and ordinal ideals
Fact 2.10
Fact 2.11

### v2

@@ 1 Introduction
Thus, by the (ZFC) results of Section 4, the poset ${\mathbb{P}}(\omega^{\delta})$ either preserves $\omega_{1}$ and forces $|{\mathfrak{c}}|=|{\mathfrak{h}}|$, or collapses (at least) $\omega_{2}$ to $\omega$. So, it is natural to ask whether ${\mathbb{P}}(\omega^{\delta})$ collapses more cardinals and a more ambitious task is to characterize the algebra $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))$. Since $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|=|P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}|\leq 2^{|\delta|}$, the extreme case is when ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega_{1}$, in cases (A) and (B), or collapses $2^{|\delta|}$ to $\omega$, in other cases. So, in Section 5 we consider that phenomenon of “maximal collapse”. For example, in cases (D) and (E) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$, if $2^{\mathop{\rm cf}\nolimits(\delta)}=2^{|\delta|}$ and ($2^{<\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)$ or $2^{\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)^{+}$). Regarding case (A) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$, for all $n\in{\mathbb{N}}$, whenever ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$, or $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and collapses $2^{|\delta|}$ to $\omega_{1}$.
@@ 2 Preliminaries
If ${\mathbb{P}}=\langle P,\leq\rangle$ is a preorder, the elements $p$ and $q$ of $P$ are incompatible, we write $p\perp q$, iff there is no $r\in P$ such that $r\leq p,q$. A set $A\subset P$ is an antichain iff $p\perp q$, for different $p,q\in A$. An element $p$ of $P$ is an atom iff each two elements $q,r\leq p$ are compatible. ${\mathbb{P}}$ is called: atomless iff it has no atoms; homogeneous iff it has a largest element and ${\mathbb{P}}\cong(\cdot,p]$, for each $p\in{\mathbb{P}}$. A set $D\subset P$ is called dense iff for each $p\in P$ there is $q\in D$ such that $q\leq p$; $D$ is called open iff $q\leq p\in D$ implies $q\in D$. If $\kappa$ is a cardinal, ${\mathbb{P}}$ is called $\kappa$-closed iff for each $\gamma<\kappa$ each sequence $\langle p_{\alpha}:\alpha<\gamma\rangle$ in $P$, such that $\alpha<\beta\Rightarrow p_{\beta}\leq p_{\alpha}$, has a lower bound in $P$. $\omega_{1}$-closed preorders are called $\sigma$-closed.
A preorder ${\mathbb{P}}=\langle P,\leq\rangle$ is called separative iff for each $p,q\in P$ satisfying $p\not\leq q$ there is $r\leq p$ such that $r\perp q$. The separative modification of ${\mathbb{P}}$ is the separative preorder $\mathop{\rm sm}\nolimits({\mathbb{P}})=\langle P,\leq^{*}\rangle$, where $p\leq^{*}q\Leftrightarrow\forall r\leq p\;\exists s\leq r\;s\leq q$. The separative quotient of ${\mathbb{P}}$ is the separative partial order $\mathop{\rm sq}\nolimits({\mathbb{P}})=\mathop{\rm asq}\nolimits(\mathop{\rm sm}\nolimits({\mathbb{P}}))$, where $\mathop{\rm asq}\nolimits({\mathbb{P}})$ denotes the antisymmetric quotient of a preorder ${\mathbb{P}}$. Preorders ${\mathbb{P}}$ and ${\mathbb{Q}}$ are called forcing equivalent, in notation ${\mathbb{P}}\equiv_{forc}{\mathbb{Q}}$, iff they produce the same generic extensions.
@@
(Folklore) If ${\mathbb{P}}$ and ${\mathbb{P}}_{i}$, for $i\in I$, are preorders, then
(a) ${\mathbb{P}}\equiv_{forc}\mathop{\rm sm}\nolimits({\mathbb{P}})\equiv_{forc}\mathop{\rm sq}\nolimits({\mathbb{P}})\equiv_{forc}\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))$;
(b) If ${\mathbb{P}}_{i}$, $i\in I$, are $\kappa$-closed preorders, then $\prod_{i\in I}{\mathbb{P}}_{i}$ is $\kappa$-closed;
(c) $\mathop{\rm sm}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})=\prod_{i\in I}\mathop{\rm sm}\nolimits{\mathbb{P}}_{i}$ and $\mathop{\rm sq}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})\cong\prod_{i\in I}\mathop{\rm sq}\nolimits{\mathbb{P}}_{i}$;
(d) If $X$ is an infinite set, ${\mathcal{I}}\subset P(X)$ is an ideal containing $[X]^{<\omega}$, ${\mathcal{I}}^{+}=P(X)\setminus{\mathcal{I}}$ is the corresponding co-ideal and $A\subset_{\mathcal{I}}B$ iff $A\setminus B\in{\mathcal{I}}$, for $A,B\in{\mathcal{I}}^{+}$, then $\mathop{\rm sm}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle$ $=\langle{\mathcal{I}}^{+},\subset_{\mathcal{I}}\rangle$ and $\mathop{\rm sq}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle=(P(X)/{\mathcal{I}})^{+}$.
If $\langle{\mathbb{P}},\leq_{\mathbb{P}},1_{\mathbb{P}}\rangle$ and $\langle{\mathbb{Q}},\leq_{\mathbb{Q}},1_{\mathbb{Q}}\rangle$ are preorders, then a mapping $f:{\mathbb{P}}\rightarrow{\mathbb{Q}}$ is a complete embedding, in notation $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$ iff
@@
Fact 2.2
If $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$, then ${\mathbb{Q}}$ is forcing equivalent to the two step iteration ${\mathbb{P}}\ast\pi$, where $\pi$ is a ${\mathbb{P}}$-name for a preorder. In addition, $F:\mathop{\rm asq}\nolimits({\mathbb{P}})\hookrightarrow_{c}\mathop{\rm asq}\nolimits({\mathbb{Q}})$, where $F([p]_{\mathop{\rm asq}\nolimits({\mathbb{P}})})=[f(p)]_{\mathop{\rm asq}\nolimits({\mathbb{Q}})}$, for $p\in{\mathbb{P}}$.
@@
If ${\mathbb{P}}_{i}\hookrightarrow_{c}{\mathbb{P}}_{i}^{\prime}$, for $i\in I$, then $\prod_{i\in I}{\mathbb{P}}_{i}\hookrightarrow_{c}\prod_{i\in I}{\mathbb{P}}_{i}^{\prime}$.
@@
If $V\models\mathrm{GCH}$ and, in $V$, $\kappa>\omega$ is a regular cardinal and ${\mathbb{P}}=\mathop{\mathrm{Fn}}\nolimits(\kappa,2)$ is the Cohen forcing, then in $V[G]$ we have ${\mathfrak{c}}=\kappa$. If, in $V$, $\mu$ and $\theta$ are infinite cardinals, then $(\theta^{\mu})^{V[G]}=(\max(\kappa,\theta)^{\mu})^{V}$ (see [7], p. 246).
@@
Let $\kappa\geq 2$ and $\lambda\geq\omega$ be cardinals. The collapsing algebra $\mathop{\rm Col}\nolimits(\lambda,\kappa)$ is the Boolean completion of the reversed tree $\langle{}^{<\lambda}\kappa,\supset\rangle$. The following fact is folklore; see [13], Thm 3.1 for a proof.
Fact 2.5
Let $\lambda\geq\omega$ be a regular cardinal and ${\mathbb{P}}$ a separative $\lambda$-closed preorder of size $\kappa=\kappa^{<\lambda}$.
(a) If $\kappa>\lambda$ and $1_{\mathbb{P}}\Vdash|\check{\kappa}|=\check{\lambda}$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\cong\mathop{\rm Col}\nolimits(\lambda,\kappa)$;
(b) If $\kappa=\lambda$ and ${\mathbb{P}}$ is atomless, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\cong\mathop{\rm Col}\nolimits(\lambda,\lambda)$.
For a cardinal $\kappa\geq\omega$ the partial order $(P(\kappa)/[\kappa]^{<\kappa})^{+}$ is denoted by ${\mathcal{P}}_{\kappa}$; thus, ${\mathcal{P}}_{\omega}=(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. If $\subset^{*}$ is the preorder on the set $[\kappa]^{\kappa}$ given by $A\subset^{*}B$ iff $|A\setminus B|<\kappa$, then, clearly, $\mathop{\rm sm}\nolimits\langle[\kappa]^{\kappa},\subset\rangle=\langle[\kappa]^{\kappa},\subset^{*}\rangle$ and $\mathop{\rm sq}\nolimits\langle[\kappa]^{\kappa},\subset\rangle={\mathcal{P}}_{\kappa}$.
@@
If $\kappa$ is an infinite cardinal, then for the partial order ${\mathcal{P}}_{\kappa}$ we have
(a) ${\mathcal{P}}_{\kappa}$ is atomless, homogeneous, of size $2^{\kappa}$ and $\kappa^{++}\leq\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})\leq(2^{\kappa})^{+}$;
(b) Forcing by ${\mathcal{P}}_{\omega}$ collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$ (Balcar, Pelant and Simon, [1]);
(c) If $\kappa>\mathop{\rm cf}\nolimits(\kappa)=\omega$, then $\mathop{\rm Col}\nolimits(\omega_{1},\kappa^{\omega})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$ (Kojman, Shelah [6]);
(d) If $\kappa>2^{\mathop{\rm cf}\nolimits(\kappa)}>\mathop{\rm cf}\nolimits(\kappa)>\omega$, then $\mathop{\rm Col}\nolimits(\omega,\kappa^{+})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$ (Shelah, [20]);
(e) If $\kappa=\mathop{\rm cf}\nolimits(\kappa)>\omega$, then ${\mathcal{P}}_{\kappa}$ collapses each $\lambda<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ to $\omega$ (Shelah, [20]).
Linear orders. Indecomposable ordinals